OpenStudy (anonymous):

What is the y-intercept of y = 0.50(1.06)x?

OpenStudy (experimentx):

the line passes through, (0,0), therefore, 0 is y intercept

Directrix (directrix):

Is y = 0.50(1.06)x supposed to be y = 0.50(1.06)^x

OpenStudy (anonymous):

yep

Directrix (directrix):

y = 0.50(1.06)^x The y-intercept appears when x = 0. y = 0.50 [ (1.06)^0] y = .5 * 1 = .5